# Regularization in Data-driven Predictive Control: A Convex Relaxation Perspective

### v1

@@ 5 Hybrid preprocessing and regularization of data for predictive control
where $\tilde{H}_{y},\tilde{H}$ denote $\textnormal{col}(\tilde{Y}_{\textnormal{P}},\tilde{Y}_{\textnormal{F}})$,and $\textnormal{col}(U_{\textnormal{P}},\tilde{Y}_{\textnormal{P}},U_{\textnormal{F}},\tilde{Y}_{\textnormal{F}})$, respectively. Compared with (10), the decision variables in (34) become $\tilde{H}_{y}$ and $K$, as we only denoise the output trajectory. Still, problem (34) is difficult to solve due to the interplay between (34a) to (34c). Without (34b), (34c), it is a structured low-rank approximation (SLRA) problem [46, 47]. We here adapt an iterative SLRA algorithm in [47] to get the approximation solution to (34).
@@ 5.1 Sequential optimization
Step 1: Low-rank approximation. Since the input data $u_{\textnormal{d}}$ has no noise and satisfies the persistent excitation, we have $\textrm{rank}(H_{u})=mL$. In contrast, the output data $y_{\textnormal{d}}$ contains “variance” noise and “bias” error, and thus the rank of the raw data $H$ (i.e., $\textnormal{col}(H_{u},H_{y})$) is larger than $mL+n$. We first consider the constraint (34a) that approximates $H$ with a low-rank matrix, which is
@@
$H_{y_{1}}:=H_{y}\Pi_{2}+\sum_{i=1}^{n}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}.$ (36)
Since we have $\mathrm{rank}(H_{u})=mL$, the key insight for Proposition 5 is to ensure the part of $H_{y}$ in the null space of $H_{u}$ has rank $n$. Specifically, we first divide $H_{y}$ into two parts that are $H_{y}\Pi_{2}$ in the row space of $H_{u}$ and $H_{y}(I-\Pi_{2})$ in the null space of $H_{u}$. We then perform an SVD of $H_{y}(I-\Pi_{2})$ to estimate its rank-$n$ approximation, and finally combine it with the component of $H_{y}$ in the row space of $H_{u}$ as in (36).
@@
Step 2: Hankel structure: We then project $H_{y_{1}}$ to a Hankel matrix set via averaging skew-diagonal elements [47] and represent the projector and the resulting Hankel matrix as $\Pi_{\textnormal{H}}$ and $H_{y_{2}}$.
@@
We repeat Steps 1 - 3 iteratively, and this process is listed in Algorithm 1. The resulting matrix $H_{y}^{*}$ from Algorithm 1 is partitioned as $\textnormal{col}(Y_{\textnormal{P}}^{*},Y_{\textnormal{F}}^{*})$, and we form a new Hankel matrix $H_{\textnormal{op}}^{*}=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}}^{*},U_{\textnormal{F}},Y_{\textnormal{F}}^{*})$. We call $H_{\textnormal{op}}^{*}$ an approximated optimal linear representation, as it directly tackles the inner identification problem without relaxation.
Algorithm 1 Iterative Construction for Approximatedly Optimal Trajectory Library
@@ 5.2 Data-driven predictive control with approximated linear data-driven representation
We use the approximated optimal linear model $H_{\textnormal{op}}^{*}$ from Algorithm 1 as the predictor in data-driven predictive control, which leads to
@@
where the $l_{1}$-norm regularizer is to implicitly select the trajectory in $H_{\textnormal{op}}^{*}$. We call this formulation (38) as Optimality-based DDPC (O-DDPC).
Note that utilizing SVD and projecting the resulting matrix to the set of Hankel matrices iteratively has been well-studied; see, e.g., [46, 47]. The convergence for the iterative algorithm using Hankel projection only is discussed in [48]. However, in Algorithm 1, we require the extra separation step to divide $H_{y}$ into the row and null spaces of $H_{u}$ and the causality projection. While our numerical simulations indicate that Algorithm 1 converges in practice, establishing a formal theoretical proof remains an open problem, which we leave for future work.
Remark 3 (LTI system characterizations).
The constraints (34a)-(34c) are necessary for the Hankel matrix $H$ to come from an LTI system, but they are not sufficient. As discussed in [1], the parameter $K$ needs to satisfy additional structure requirements to ensure the existence of the corresponding matrix parameters $A,B,C,D$ in (1). When data is collected from nonlinear systems, making the data more structured may introduce a larger bias error. Our proposed O-DDPC in (38) preserves the linear, low-rank structure, and causality without requiring the predictor $H_{\textnormal{op}}^{*}$ to exactly correspond to an LTI system. This hybrid preprocessing and regularization imposes more system structure but also gives freedom may allow (38) to select model complexity implicitly, thus improving empirical control performance.
@@ 6 Numerical experiments
We compare the numerical performance [^1] Our code is available at https://github.com/soc-ucsd/Convex-Approximation-for-DeePC. for the DDPC variants in Sections 4 and 5, including: 1) Linearity-based DDPC (23) (L-DDPC), 2) Causality-based DDPC (29) (C-DDPC) and 3) Optimality-based DDPC (38) (O-DDPC). We also include results from the Hankel-form SPC (SPC) in (21) and the indirect DDPC using system identification for comparison. In particular, we used the N4SID [49] for the system ID. We do not consider (33) individually, as an $l_{1}$-norm regularizer is added for all DDPC variants and the Hankel-form SPC.
We perform two sets of experiments: 1) an LTI system with noisy measurement [17] and 2) another nonlinear system with noise-free measurement [1] (corresponding to “variance” error and “bias” error, per [1]). Our results show the superior performance of O-DDPC in both systems. We also see that the indirect method performs better under the “variance” error, while the direct DDPC variants have superior performance under the “bias” error. This finding is consistent with [1].
6.1 An LTI system
Experiment setup. We first consider an LTI system from [17], which is a triple-mass-spring system with $n=8$ states, $m=2$ inputs (two stepper motors), and $p=3$ outputs (disc angles). We consider data collection with additive Gaussian measurement noises $\omega\sim\mathcal{N}(0,\sigma I)$. We utilize the data sets with various sizes and noise levels that are $T=400,600,800$ and $\sigma$ from $0.02$ to $0.1$ in increments of $0.02$. The prediction horizon and the initial sequence are chosen as $N=40$ and $T_{\textnormal{ini}}=4$, respectively. We choose $Q=I$, $R=0.1I,\mathcal{U}=[-0.7,0.7]$ and $\Gamma,\mathcal{Y}=\mathbb{R}^{3}$ unless otherwise specified.
@@ 6.1 An LTI system
Figure 4: Comparison of realized control cost of different variants of DDPC with varying measurement noises. (a), (b) and (c) illustrate the control performance of data-driven controllers with different pre-collected trajectory lengths. Across all experiments, the proposed O-DDPC (38) is comparable to the system ID approach, which provides the best control performance among all DDPC methods.
The numerical results are shown in Fig. 4. As expected, the control performance of all data-driven controllers deteriorates with increasing noise levels. Among the DDPC-based methods, the proposed O-DDPC (38) achieves the best realized control costs and exhibits the strongest noise robustness. This improvement may come from its iterative low-rank approximation, causality projection, and Hankel structure projection (see Algorithm 1), which together act as an effective noise filter. This reduces the influence of variance noises in LTI systems.
@@
O-DDPC 300.77 ($\mathbf{8.48\%}$) 296.59 ($\mathbf{6.98\%}$) 296.31 ($\mathbf{6.88\%}$)
@@
Table 1 lists the realized control costs at noise level $\sigma=0.1$ for different pre-collected trajectory lengths and control approaches. We also report their percentage increases relative to the ground-truth cost computed with noise-free data from (6). The results show a clear performance ordering: $\texttt{L-DDPC}>\texttt{SPC}>\texttt{C-DDPC}>\texttt{O-DDPC}>\textrm{System ID}$. For the LTI system with noisy data, the inner problem in (7) enforces LTI structure in the data-driven representation, which in turn improves noise rejection in the outer predictive control problem (7). Consequently, methods that impose more constraints on the inner problem tend to yield more structured representations and better performance. Notably, the increase in realized control cost for O-DDPC is approximately $7\%$, significantly outperforming the other DDPC variants.
@@ 6.2 Nonlinear system
We here consider a nonlinear Lotka-Volterra system [1]
$\dot{x}=\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}ax_{1}-bx_{1}x_{2}\\ dx_{1}x_{2}-cx_{2}+u\end{bmatrix}$, (39)
where $x_{1},x_{2}$ denote prey and predator populations and $u$ is the control input. We used $a=c=0.5,b=0.025,d=0.005$ in our experiments. We first linearize the nonlinear system (39) around the equilibrium $(\bar{u},\bar{x}_{1},\bar{x_{2}})=(0,\frac{c}{d},\frac{a}{b})$. Then, after discretization, we obtain a linear system as
$\hat{x}(k+1)$ $=f_{\text{linear}}(\hat{x}(k),\hat{u}(k))$
$=\begin{bmatrix}\hat{x}_{1}(k)+\Delta t(-b\bar{x}_{1}\hat{x}_{2}(k))\\ \hat{x}_{2}(k)+\Delta t(d\bar{x}_{2}\hat{x}_{1}(k)+\hat{u}(k))\end{bmatrix}$
where $\Delta t=0.1$ is the time step for discretization. We discretize the nonlinear system in the error state space
$\hat{x}(k+1)$ $=f_{\text{nonlinear}}(\hat{x}(k),\hat{u}(k))$
$=\begin{bmatrix}\hat{x}_{1}(k)+\Delta t(a(\hat{x}_{1}(k)+\bar{x_{1}})\\ -b(\hat{x}_{1}(k)+\bar{x_{1}})(\hat{x}_{2}(k)+\bar{x_{2}}))\\ \hat{x}_{2}(k)+\Delta t(d(\hat{x}_{1}(k)+\bar{x_{1}})(\hat{x}_{2}(k)+\bar{x_{2}})\\ -c(\hat{x}_{2}(k)+\bar{x_{2}})+\hat{u}(k))\end{bmatrix}$.
We then construct systems with various nonlinearity by interpolating between $f_{\textrm{linear}}$ and $f_{\textrm{nonlinear}}$ that is
@@
$\quad+(1-\epsilon)\cdot f_{\text{nonlinear}}(\hat{x}(k),\hat{u}(k)).$
The length of the pre-collected trajectory is $T=400$. The prediction horizon and initial sequence are set as $N=60$ and $T_{\textnormal{ini}}=4$, respectively. We choose $Q=I,R=0.5I$ and $\mathcal{U}=[-20,20]$. The parameters for each controller are the same as those in Section 6.1.
Comparison of direct/indirect methods. We compare the realized control costs for the indirect system ID approach and different DDPC variants on systems with varying degrees of nonlinearity. Model orders are chosen to be 8 and 4 for O-DDPC in Algorithm 1 and N4SID, respectively. Similar to Section 6.1, we average the realized control costs over 100 pre-collected trajectories. We note that the identified model from N4SID is often ill-conditioned when the nonlinearity is high, which caused numerical issues in solving (2) in some of our experiments. We discard these ill-conditioned solutions when computing the average performance for the indirect system ID approach.
Both direct and indirect approaches perform well when the nonlinearity is low (i.e., $\epsilon\in[0.8,1])$, as shown in Fig. 5. However, the cost for the indirect method significantly increases with higher nonlinearity, while the performance of direct methods remains relatively consistent. The superior performance of direct data-driven methods is consistent with experimental observations from [1]. The indirect system ID method projects the noisy data on a fixed linear model, which induces “bias” error due to selecting a wrong model class; on the other hand, the complexity of the LTI system is regularized but not specified in direct methods, which provides more flexibility and leads to superior performance for controlling nonlinear systems in our experiments.
@@
Comparison of DDPC variants. We then compare the performance of different DDPC variants. The results are shown in Fig. 6. The C-DDPC and O-DDPC with more structured data-driven representation outperform L-DDPC and SPC which only relax or tackle the linearity requirement without considering the causality and the Hankel structure. Furthermore, among all direct data-driven approaches, O-DDPC performs the best for both the nonlinear system (bias error) and the LTI system with measurement noise (variance noise). These numerical results suggest that we might obtain additional benefits when employing appropriate techniques from system ID to pre-process the trajectory library of nonlinear systems.
Figure 6: Comparison of realized control cost of different DDPC variants for systems with varying nonlinearity. Across all experiments, the proposed O-DDPC (38) provides the best control performance among all DDPC methods.

### v2

@@ 5 Hybrid preprocessing and regularization of data for predictive control
where $\tilde{H}_{y},\tilde{H}$ denote $\textnormal{col}(\tilde{Y}_{\textnormal{P}},\tilde{Y}_{\textnormal{F}})$, and $\textnormal{col}(U_{\textnormal{P}},\tilde{Y}_{\textnormal{P}},U_{\textnormal{F}},\tilde{Y}_{\textnormal{F}})$, respectively. Compared with (10), the decision variables in (34) become $\tilde{H}_{y}$ and $K$, as we only denoise the output trajectory. Still, problem (34) is difficult to solve due to the interplay between (34a) to (34c). Without (34b), (34c), it is a structured low-rank approximation (SLRA) problem [markovsky2008structured, yin2021low]. We here adapt an iterative SLRA algorithm in [yin2021low] to get the approximation solution to (34).
@@ 5.1 Sequential optimization
Step 1: Low-rank approximation. Since the input data $u_{\textnormal{d}}$ has no noise and satisfies the persistent excitation, we have $\textrm{rank}(H_{u})=mL$. In contrast, the output data $y_{\textnormal{d}}$ may contain “variance” noise and “bias” error. Thus, the rank of the raw data $H$ (i.e., $\textnormal{col}(H_{u},H_{y})$) is larger than $mL+n$. We first consider the constraint (34a) that approximates $H$ with a low-rank matrix, which is
@@
$H_{y_{1}}:=H_{y}\Pi_{2}+\sum_{i=1}^{n}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}.\vskip-5.69054pt$ (36)
Since we have $\mathrm{rank}(H_{u})=mL$, the key insight for Proposition 5 is to ensure the part of $H_{y}$ in the null space of $H_{u}$ has rank $n$. Specifically, we first divide $H_{y}$ into two parts that are $H_{y}\Pi_{2}$ in the row space of $H_{u}$ and $H_{y}(I-\Pi_{2})$ in the null space of $H_{u}$. We then perform an SVD of $H_{y}(I-\Pi_{2})$ to estimate its rank-$n$ approximation, and finally combine it with the component of $H_{y}$ in the row space of $H_{u}$ as in (36). A detailed proof is provided in Appendix B.5.
@@
Step 2: Hankel structure: We then project $H_{y_{1}}$ to a Hankel matrix set via averaging skew-diagonal elements [yin2021low] and represent the projector and the resulting Hankel matrix as $\Pi_{\textnormal{H}}$ and $H_{y_{2}}$.
@@
We repeat Steps 131\!-\!3 iteratively as listed in Algorithm 1. The resulting matrix $H_{y}^{*}$ from Algorithm 1 is partitioned as $\textnormal{col}(Y_{\textnormal{P}}^{*},Y_{\textnormal{F}}^{*})$, and we form a new Hankel matrix $H_{\textnormal{op}}^{*}=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}}^{*},U_{\textnormal{F}},Y_{\textnormal{F}}^{*})$. We call $H_{\textnormal{op}}^{*}$ an approximated optimal linear representation, as it directly tackles the inner identification problem without relaxation.
Algorithm 1 Iterative Construction for Approximately Optimal Trajectory Library
@@ 5.2 Data-driven predictive control with approximated linear data-driven representation
We use the approximately optimal linear model $H_{\textnormal{op}}^{*}$ from Algorithm 1 as the predictor in data-driven predictive control, which leads to
@@
where the $l_{1}$-norm regularizer is used to implicitly select the trajectory in $H_{\textnormal{op}}^{*}$. We call this formulation (38) as Approximation-based DDPC (A-DDPC). We note that $H_{\textnormal{op}}^{*}$ only needs to be computed once via the offline collected data, and the initial trajectories $u_{\textnormal{ini}},y_{\textnormal{ini}}$ are updated online at each step.
The SLRA [markovsky2008structured, yin2021low], which utilizes SVD and projection onto the set of Hankel matrices iteratively, has been well-studied. Sufficient conditions for its convergence guarantee are discussed in [andersson2013alternating]. However, these results cannot be directly applied to Algorithm 1 as we require extra separation steps to divide $H_{y}$ into the row and null spaces of $H_{u}$ and the causality projection. In numerical implementation, Algorithm 1 is applied one time offline and was observed to converge in all our experiments with various pre-collected trajectories. Establishing a formal theoretical proof and analyzing its convergence rate remain open, which we leave for future work.
Remark 4 (LTI systems and beyond).
Constraints (34a)-(34c) are necessary for the Hankel matrix $H$ to come from an LTI system, but they are not sufficient. As discussed in [dorfler2022bridging], the parameter $K$ needs to satisfy additional structure requirements to ensure the existence of the corresponding matrix parameters $A,B,C,D$ in (1). For nonlinear systems, making the data matrix $H$ more structured, as in the LTI setting, may introduce a larger bias error. Our proposed A-DDPC in (38) preserves the linear, low-rank structure, and causality without requiring the predictor $H_{\textnormal{op}}^{*}$ to exactly correspond to an LTI system. This hybrid preprocessing and regularization imposes certain system structure but also gives flexibility to allow (38) to select model complexity implicitly, thus improving empirical control performance.
Remark 5 (Complexity of DDPC variants).
L-DDPC (23), C-DDPC (29), and A-DDPC (38) have the same optimization form and thus share similar online computational complexity. Their main difference lies in the construction of trajectory libraries and the choice of regularizers. Using the equality constraints, we can eliminate $u,y,\sigma_{y}$, so that the only decision variable is $g$, whose dimension is $T-T_{\textnormal{ini}}-N+1$. For standard LTI systems, the length of the pre-collected trajectory to ensure the persistent exciting of the input is $(m+1)(T_{\textnormal{ini}}+N+n)-1$, and we can choose $T_{\textnormal{ini}}=n$. In this case, $g$ has dimension $m(N+2n)+n$, which is $2mn+n$ larger than that in classical MPC and is thus of comparable order. For systems beyond LTI settings, a longer pre-collected trajectory is generally needed to capture richer system behavior. In our experiments, the computation time for systems beyond LTI settings is on the order of $10^{-2}\,\mathrm{s}$, which remains suitable for real-time implementation.
@@ 6 Numerical experiments
We compare the numerical performance [^1] Our code is available at https://github.com/soc-ucsd/Convex-Approximation-for-DeePC. for the DDPC variants in Sections 4 and 5, including: 1) L-DDPC (23), 2) C-DDPC (29) and 3) A-DDPC (38). We also include results from the Equality-form SPC (SPC) in (21) and the indirect DDPC using system identification for comparison. In particular, we used the N4SID [van1994n4sid] for the system ID. We do not consider (33) individually, as an $l_{1}$-norm regularizer is added for all DDPC variants.
We perform two sets of experiments in this section: 1) an LTI system with noisy measurement [fiedler2021relationship] (“variance” error) and 2) another nonlinear system with noise-free measurement [dorfler2022bridging] (“bias” error). An additional example of controlling a two-wheeled robot using DDPC variants is provided in Appendix C. We see that the indirect method performs better on the non-deterministic LTI system under the “variance” error, while the direct DDPC variants have superior performance on the nonlinear system under the “bias” error. This finding is consistent with [dorfler2022bridging]. Furthermore, our results show the superior performance of A-DDPC in both systems. For hyperparameter selection, we perform a grid search for $\lambda_{g}$ and $\lambda_{1}$ over a broad range for the non-deterministic LTI system: we use the realized cost as the evaluation metric and choose a region that yields relatively good performance. The selected parameters are further tuned in the nonlinear example to enhance numerical stability.
6.1 Non-deterministic LTI system
Experiment setup. We first consider an LTI system from [fiedler2021relationship], which is a triple-mass-spring system with $n=8$ states, $m=2$ inputs (two stepper motors), and $p=3$ outputs (disc angles). We consider data collection with additive Gaussian measurement noises $\omega\sim\mathcal{N}(0,\sigma I)$. We utilize the data sets with various sizes and noise levels that are $T=400,600,800$ and $\sigma$ from $0.02$ to $0.1$ in increments of $0.02$. The prediction horizon and the initial sequence are chosen as $N=40$ and $T_{\textnormal{ini}}=4$, respectively. We choose $Q=I$, $R=0.1I,\mathcal{U}=[-0.7,0.7]$ and $\Gamma,\mathcal{Y}=\mathbb{R}^{3}$ unless otherwise specified.
@@ 6.1 Non-deterministic LTI system
Figure 4: Comparison of realized control cost of different variants of DDPC with varying measurement noises. (a), (b) and (c) illustrate the control performance of data-driven controllers with different pre-collected trajectory lengths. Across all experiments, the proposed A-DDPC (38) is comparable to the system ID approach, which provides the best performance among all DDPC methods.
The numerical results are shown in Fig. 4. As expected, the control performance of all data-driven controllers deteriorates with increasing noise levels. Among the DDPC-based methods, the proposed A-DDPC (38) achieves the best realized control costs and exhibits the strongest noise robustness. This improvement may come from its iterative low-rank approximation, causality projection, and Hankel structure projection (see Algorithm 1), which together act as an effective noise filter. This reduces the influence of variance noises in LTI systems.
@@
A-DDPC 300.77 ($\mathbf{8.48\%}$) 296.59 ($\mathbf{6.98\%}$) 296.31 ($\mathbf{6.88\%}$)
@@
Table 1 lists the realized control costs at noise level $\sigma=0.1$ for different pre-collected trajectory lengths and control approaches. We also report their percentage increases relative to the ground-truth cost computed with noise-free data from (6). The results show a clear performance ordering: $\texttt{L-DDPC}>\texttt{SPC}>\texttt{C-DDPC}>\texttt{A-DDPC}>\textrm{System ID}$. For the LTI system with noisy data, the inner problem in (7) enforces LTI structure in the data-driven representation, which in turn improves noise rejection in the outer predictive control problem (7). Consequently, methods that impose more constraints on the inner problem tend to yield more structured representations and better performance. Notably, the increase in realized control cost for A-DDPC is approximately $7\%$, significantly outperforming the other DDPC variants.
Fig. 5 shows typical trajectories for all methods under different noise levels. In this case, the open-loop trajectories for all approaches remain close to the optimal trajectory obtained by MPC with the accurate system model up to $1\,\mathrm{s}$. After that, $\texttt{L-DDPC},\texttt{SPC}$ and C-DDPC exhibit larger oscillations and deviations from the optimal trajectory, whereas the proposed A-DDPC and the sequential system identification and control are closer to the optimal one with reduced oscillations. Furthermore, across different noise levels, L-DDPC, SPC and C-DDPC show relatively large variability while A-DDPC and the system identification approach are less sensitive to noise.
(a) L-DDPC
(b) SPC
(c) C-DDPC
(d) A-DDPC
(e) System ID
Figure 5: Output trajectories (the angle of disc 2) of different approaches. The blue trajectory denotes the optimal trajectory using MPC with the true system model. The orange trajectories from dark to light represent different DDPC variants under noise levels of $0.02$, $0.06$ and $0.1$, respectively.
@@ 6.2 Nonlinear system
We here consider a nonlinear Lotka-Volterra system [dorfler2022bridging]
[…]
where $x_{1},x_{2}$ denote prey and predator populations and $u$ is the control input. We use $a=c=0.5,b=0.025,d=0.005$ and the discretization time step $\Delta t=0.1$ in our experiments. We denote the discretized form of the nonlinear system (39) and its linearization around the equilibrium $(\bar{u},\bar{x}_{1},\bar{x_{2}})=(0,\frac{c}{d},\frac{a}{b})$ in the error state space as $\hat{x}(k+1)=f_{\text{nonlinear}}(\hat{x}(k),\hat{u}(k))$ and $\hat{x}(k+1)=f_{\text{linear}}(\hat{x}(k),\hat{u}(k))$, respectively.
We then construct systems with various nonlinearity, which gradually depart from the linear regime and the violation of the LTI assumption becomes more significant. We aim to evaluate how the performance of different DDPC approaches evolves as the level of model mismatch increases. Specifically, we interpolate between $f_{\textrm{nonlinear}}$ and $f_{\textrm{linear}}$ that is
@@
$\quad+(1-\epsilon)\cdot f_{\text{nonlinear}}(\hat{x}(k),\hat{u}(k)),$
and further collect the input-output trajectories of these systems with different $\epsilon$ ranging from 0 to $1$ to formulate the associated trajectory libraries. The length of the pre-collected trajectory is $T=400$. The prediction horizon and initial sequence are set as $N=60$ and $T_{\textnormal{ini}}=4$. We choose $Q=I,R=0.5I$ and $\mathcal{U}=[-20,20]$.
Comparison of direct/indirect methods. We compare the realized control costs for the indirect system ID approach and different DDPC variants on systems with varying degrees of nonlinearity. Model orders are chosen to be 8 and 4 for A-DDPC in Algorithm 1 and N4SID, respectively, as they yield relatively satisfactory performance among various orders from testing. Similar to Section 6.1, we average the realized control costs over 100 pre-collected trajectories. We note that the identified model from N4SID is often ill-conditioned when the nonlinearity is high, which caused numerical issues in solving (2) in some of our experiments. We discard these ill-conditioned solutions when computing the average performance for the indirect system ID approach.
Both direct and indirect approaches perform well when the nonlinearity is low (i.e., $\epsilon\in[0.8,1])$, as shown in Fig. 6. However, the cost for the indirect method significantly increases with higher nonlinearity, while the performance of direct methods remains relatively consistent. The superior performance of direct data-driven methods is consistent with experimental observations from [dorfler2022bridging]. The indirect system ID method projects the noisy data on a structured linear model parameterized by the matrices $(A,B,C,D)$ (see Remark 4), which induces “bias” error due to selecting a wrong model class. On the other hand, the complexity of the LTI system is regularized but not specified in direct methods, such as L-DDPC and C-DDPC, where suitable choices of the weighting parameters provide more flexibility. For the A-DDPC, the model order $n$ in (34a) is tunable and can be chosen larger than the state dimension, and it does not require identification of an explicit parametric LTI model, which gives an improved capacity to capture the richer behavior of the nonlinear system. These features lead to superior performance of L-DDPC, C-DDPC, and A-DDPC for controlling nonlinear systems in our experiments.
@@
Comparison of DDPC variants. We then compare the performance of different DDPC variants. The results are shown in Fig. 7. It is obvious that realization costs for all approaches become higher with the increase of nonlinearity. The C-DDPC and A-DDPC with more structured data-driven representation outperform L-DDPC and SPC which only relax or tackle the linearity requirement without considering the causality and the Hankel structure. Furthermore, among all direct data-driven approaches, A-DDPC performs the best for both the nonlinear system (bias error) and the LTI system with measurement noise (variance noise). These numerical results suggest that we might obtain additional benefits when employing appropriate techniques from system ID to pre-process the trajectory library of nonlinear systems. Quantitatively analyzing the effect of nonlinearity is an interesting direction for future work, and the distance to the class of nonlinear systems that admit an accurate Koopman linear model [shang2026existence] may serve as a useful metric.
Figure 7: Comparison of realized control cost of different DDPC variants for systems with varying nonlinearity. Across all experiments, the proposed A-DDPC (38) provides the best control performance among all DDPC methods.